1) The first number is 5 more than the 2nd. Three times the first increased by twice the second results in 30. Find the numbers.

Found 2 solutions by Boreal, stanbon:
Answer by Boreal(15235)

x is first number
x-5 is the second
3x+2(x-5)=30
3x+2x-10=30
5x=40
x=8
x-5=3
the numbers are 8 and 3
8 is 5 more than 3
3*8=24 and increased by 2(3) makes 30.

Answer by stanbon(75887)

The first number is 5 more than the 2nd. Three times the first increased by twice the second results in 30. Find the numbers.
------------------------
f = s+5
3f + 2s = 30
------------------------------
Substitute for "f" and solve for "s"::
3(s+5) + 2s = 30
---------------------------
3s + 15 + 2s = 30
5s = 15
s = 3 (the 2nd number)
f = s+5 = 8 (the 1st number)
--------------------------------
Cheers,
Stan H.
----------------